an earthquake was felt throughout a circular area of 1,808.64 square miles. what is the radius of the circular area?
a- 24 miles
b- 576 miles
c- 12 miles
d- 20 miles

Answer by Earlsdon(6098)

The area of a circle is given by: A = (pi)(r^2)
The problem gives the area as: 1,808.64 sq. mi.
So, (pi)(r^2) = 1,808.64 Solve for r. Divide both sides by pi (3.14)
r^2 = 1.808.64/3.14
r^2 = 576 Take the square root of both sides.
r = 24 Miles.